A company had total revenues of $104 million, operating profit margin of 24%, and depreciation and amortization expense of $9 million over the trailing twelve months. The company currently has $49 million in total debt and $10 million in cash and cash equivalents. If the company's market capitalization (market value of its equity) is $593 million, what is its EV/EBITDA ratio? Round to one decimal place.
Answer #1

EBITDA=104*24%+9=33.96 million

what is its EV/EBITDA ratio

=(593+49-10)/33.96

=18.6

the above is answer..